OpenStudy (anonymous):

Solve using matrices. 4x + 5y = 40 x – y = 1

OpenStudy (unklerhaukus):

\[\left[\begin{array}{cc|c}4&5&40\\1&-1&1\end{array}\right]\]

OpenStudy (unklerhaukus):

\[R_1\to R_1-3R_2\]\[\left[\begin{array}{cc|c}1&8&37\\1&-1&1\end{array}\right]\] \[R_2\to R_2-R_1\]\[\left[\begin{array}{cc|c}1&8&37\\0&-9&-36\end{array}\right]\]

OpenStudy (unklerhaukus):

can you tell me what the next row operation will be ?

OpenStudy (anonymous):

|4 5||x| |40| | || |=| | |1 -1||y| |1 | find inverse of |4 5| | | |1 -1| and then multiply both sides and find x and y